Circular No. 6989

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998dg IN MCG –3-34-008 B. Schmidt, Mount Stromlo and Siding Spring Observatories, reports

that spectrophotometry (range 380–750 nm) obtained with the Australian National University’s 2.3-m telescope at Siding Spring Observatory on Aug. 13.5 UT shows SN 1998dg (cf. IAUC 6978) to be a type-Ia supernova, ∼ 2 weeks past maximum light. A redshift was unavailable for the galaxy, but comparison of the supernova spectrum with other type-Ia supernovae suggests that SN 1998dg is at z ∼ 0.01.

SUPERNOVA 1998dh IN NGC 7541 O. Kuhn, F. Prada, and R. Costero, Instituto de Astronomía, Uni-

versidad Nacional Autonoma de México, report the following near-infrared photometric magnitudes obtained on July 28.461 UT with the 2.1-m telescope (+ CAMILA instrument) at the Observatorio Astronómico Nacional at San Pedro Martir: J =14.43 ± 0.05, H = 14.35 ± 0.05.

CCD V magnitudes from D. Hanžl, Brno, Czech Republic, using a 0.40-m reflector (GSC comparison stars): Aug. 5.933 UT, 14.0; 8.993, 14.0; 9.959, 13.95.

S/1997 U 2

Extension to the ephemeris on IAUC 6869:

1998 TT α_{2000}  δ_{2000}  Δ  r_{U}  ϵ  Δα′′ Δδ′′
Aug. 25 20^{h}51.^{m}37 −18^{o}19.^{′}7 18.855 0.114 158.^{o}6 900+ 297+
Sept. 4 20 50.01 −18 25.3 18.935 0.114 148.6 918+ 288+
14 20 48.83 −18 29.9 19.039 0.115 138.6 934+ 279+
24 20 47.91 −18 33.6 19.166 0.116 128.6 949+ 269+
Oct. 4 20 47.27 −18 36.0 19.311 0.117 118.6 963+ 260+
14 20 46.95 −18 37.2 19.471 0.118 108.7 976+ 252+
24 20 46.97 −18 37.0 19.640 0.118 98.7 989+ 244+
Nov. 3 20 47.33 −18 35.5 19.813 0.119 88.8 1001+ 236+
13 20 48.04 −18 32.6 19.986 0.120 79.0 1014+ 229+
23 20 49.07 −18 28.4 20.154 0.120 69.1 1026+ 223+
Dec. 3 20 50.41 −18 23.0 20.311 0.121 59.3 1038+ 218+
13 20 52.03 −18 16.5 20.454 0.121 49.6 1051+ 213+
1998 August 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green